OpenStudy (mtalhahassan2):

A water trough is 10 m long and a cross-section has the shape of an isosceles triangle that is 1 m across at the top and is 50 cm high. The trough is being filled with water at a rate of 0.4 m3/min. How fast will the water level rise when the water is 40 cm deep?

OpenStudy (carlyleukhardt):

Water level is rising at rate of 0.0143 m/min or 1.43 cm/min

OpenStudy (koikkara):

@MTALHAHASSAN2 I would have solved like this, Given: dv/dt = 0.4 m^3/min h=0.4 Goal: find dh/dt V=.5(lwh) = .5(10)(2h)h So V=10h^2 -------------- V=.5(lwh) dV/dt = 2(10)(h)(dh/dt) .4 =2(10)(0.4)(dh/dt) dh/dt =0.05 or 1/20 m/min